### v1

# HEAL: An Empirical Study on in mbodied gents Driven by arge Language Models

Abstract
Large language models (LLMs) are increasingly being adopted as the cognitive core of embodied agents. However, inherited hallucinations, which stem from failures to ground user instructions in the observed physical environment, can lead to navigation errors, such as searching for a refrigerator that does not exist. In this paper, we present the first systematic study of hallucinations in LLM-based embodied agents performing long-horizon tasks under sceneātask inconsistencies. Our goal is to understand to what extent hallucinations occur, what types of inconsistencies trigger them, and how current models respond. To achieve these goals, we construct a hallucination probing set by building on an existing benchmark, capable of inducing hallucination rates up to higher than base prompts. Evaluating 12 models across two simulation environments, we find that while models exhibit reasoning, they fail to resolve scene-task inconsistenciesāhighlighting fundamental limitations in handling infeasible tasks. We also provide actionable insights on ideal model behavior for each scenario, offering guidance for developing more robust and reliable planning strategies.
HEAL: An Empirical Study on in
mbodied gents Driven by arge Language Models
Trishna ChakrabortyāUdita Ghosh āXiaopan ZhangāFahim Faisal Niloy Yue Dong āJiachen Li āAmit K. Roy-Chowdhury āChengyu Song University of California, Riverside {tchak006,ughos002,xzhan006,fnilo001,jiachen.li,yue.dong}@ucr.edu csong@cs.ucr.edu āamitrc@ece.ucr.edu
1 Introduction
Recent advances in the reasoning and generalization capabilities of large language models (LLMs)Ā Chang etĀ al. (2024); Wei etĀ al. (2022) have led to their increasing adoption as the cognitive coreĀ Mai etĀ al. (2023) of embodied agentsĀ Zhang etĀ al. (2024b); Kannan etĀ al. (2024); Dorbala etĀ al. (2023), enabling these systems to interpret instructions in natural language and formulate action plans in complex environments. However, LLMs have well-known vulnerabilitiesĀ Liu etĀ al. (2024); Chakraborty etĀ al. (2024). Consequently, LLM-driven agentsĀ Xiang etĀ al. (2023); Yang etĀ al. (2025) inherit not only the vast world knowledge and reasoning capability of LLMs, but also their limitationsĀ Jiao etĀ al. (2024); Zhang etĀ al. (2025); most notably a persistent tendency to hallucinateĀ PerkoviÄ etĀ al. (2024); Sriramanan etĀ al. (2024).
While hallucination is a well-recognized limitation of LLMsĀ Tonmoy etĀ al. (2024); Rawte etĀ al. (2023a), its manifestation in embodied agents is qualitatively different. Unlike conversational systems, where hallucinations often result in factual errors or incoherent replies Ā Zhou etĀ al. (2023); Yu etĀ al. (2024a), hallucinations in embodied agents stem from a failure to ground user-provided task instructions in the observed physical environment. This misalignment can lead to consequences far more serious than a simple textual error. For example, if a robot is instructed to āput the knives in the dishwasherā but no dishwasher is present, an LLM unable to reconcile the task with the observed scene may hallucinate the existence of the dishwasher, and include it in the generated plan to follow the user instruction. This can cause the robot to place sharp utensils in an empty cabinet or try to press buttons on a bare wall, leading to physical damage, safety hazards, and wasted battery. Such behaviors highlight the need for scene-task-consistent planning in LLM-based agents.
Motivated by these limitations of current LLMsāmost of which are optimized to complete tasks under ideal conditionsāwe shift our focus to understand their failure modes. Specifically, we aim to answer the following research questions: RQ1: To what extent do LLM-based embodied agents hallucinate under sceneātask inconsistencies; what types of mismatches are more likely to trigger hallucinations; and what limitations do these failures reveal? RQ2: Does the absence of hallucination imply correct planning? What are the ideal behaviors in these scenarios for more robust planning?
Although prior works explore LLMs in embodied agentsĀ Majumdar etĀ al. (2024); Islam etĀ al. (2023) and hallucinations in QAĀ Guan etĀ al. (2024); Xu etĀ al. (2025), studies on hallucinations in embodied agents, especially for long-horizon tasks, are limited. Existing effortsĀ Zhou etĀ al. (2024) mainly study incidental cases from generic prompts, capturing only surface-level issues and overlooking the failure patterns behind hallucinations. To fill this gap, we present, to the best of our knowledge, the first empirical study that systematically exemplifies and quantifies hallucinations in long-horizon planning through controlled scene-task inconsistencies.
Because existing datasets either do not explicitly aim to elicit hallucination in long-horizon embodied tasks, or do not target embodied agent tasks, we first construct a new probing set that is more likely to cause hallucination. Constructing such a probing set faces two main challenges: (i) how to induce hallucinations and (ii) how to detect unwanted behaviors caused by hallucinations. We solve these challenges by systematically modifying base prompts from the existing LLM-based embodied agent benchmarkĀ Li etĀ al. (2024b) to introduce scene-task inconsistencies. Specifically, we modify the two core components of each base promptāthe task description and the scene informationāto create mismatches between user instructions and the observed environment (see FigureĀ 2).
With the new hallucination inducing probing set, we conduct an empirical study of popular LLMs. Overall, among our four controlled modifications, Synonymous Object Substitution yields the lowest hallucination rates, suggesting that models can recognize objects conceptually belonging to the same category but often fail to maintain naming consistency, probably due to the inherent tendency of language models to favor varied phrasing over strict lexical alignment. In contrast, the highest hallucination rates occur under Scene Task Contradiction (see FigureĀ 1), revealing the modelās inability to ground planning to perceived environments.
We also empirically study the effectiveness of mitigation strategiesĀ Peng etĀ al. (2023); Yin etĀ al. (2024). The result shows that, even with feedback-based self-correction, hallucinations in Scene Task Contradiction remain high, underscoring a fundamental inability of models to recognize the infeasibility of the task. Additionally, our small-scale experiments with vision-language models (VLMs) suggest that hallucinations are reduced when both image and text inputs are available, emphasizing the importance of cross-modal verification.
In summary, our contributions are as follows:
ā¢
We present the first study of hallucinations in LLM-based embodied agents for long-horizon tasks under sceneātask inconsistencies. Our probing set, building on the existing benchmark, elicits hallucination rates up to higher than base prompts, effectively exposing hallucinations. Our designed scenarios can be leveraged to guide robust planning.
ā¢
Our study, involving 12 models and a new hallucination probing set based on two simulation environments, reveals that among the four variants, Scene Task Contradiction induces the highest hallucination rates. The models exhibit signs of reasoning and do not blindly follow prompts; however, a prominent pattern is their inability to reject infeasible tasksāstemming from the trait that they do not know how to say ānoā.
ā¢
The absence of hallucination does not guarantee correct planning in scene-task inconsistencies. For instance, when instructed to turn on a washing machine that is not present, the models fail to reject the task and instead re-purpose available objectsāsuch as turning on the showerāresulting in unsafe actions.
2 Related Work
LLMs in Embodied AI. Use of LLMs in Embodied AI ranges from high-level planners that decompose instructions into subtasks, as shown in SayCanĀ Ahn etĀ al. (2022); to multimodal reasoners like PaLM-EĀ Driess etĀ al. (2023). LLMs can also serve as natural interaction interfaces between humans and robotsĀ Cui etĀ al. (2023). Such versatile capabilities have led to the integration of LLMs throughout the embodied AI stack, from perception processingĀ Kamath etĀ al. (2021) to decision-making frameworks combining internet knowledge with embodied groundingĀ Song etĀ al. (2023); Zawalski etĀ al. (2024); Ren etĀ al. (2023); Huang etĀ al. (2022). Vision-Language Action (VLA)Ā Jiang etĀ al. (2023); Brohan etĀ al. (2023); Kim etĀ al. (2024) models further extend these capabilities by jointly learning representations among modalities like vision, language, and physical action. Despite these advances, these systems still face grounding challenges Ā Majumdar etĀ al. (2024); Islam etĀ al. (2023).
Hallucinations in LLMs. Hallucinations in LLMs have been widely studied in text-onlyĀ Dhuliawala etĀ al. (2023); Agrawal etĀ al. (2024). Recent works extend these to multimodal LLMs Ā Bai etĀ al. (2024); Xu etĀ al. (2025); Guan etĀ al. (2024). Causes include data quality issues, weak vision encoders, and language model priors overriding visual evidence. Mitigation approaches include cross-checking Ā Yu etĀ al. (2024b), instruction-tuningĀ Liu etĀ al. (2023), self-correctionĀ Peng etĀ al. (2023), and others.
Hallucinations in Embodied AI. Recent studies Ā Li etĀ al. (2024a, b); Zhou etĀ al. (2024) have examined incidental embodied AI hallucinations from generic prompts, but these capture only surface-level issues without revealing underlying failure patterns. Other work Ā Yang etĀ al. (2024b) investigates object hallucinations using binary existence questions, but such simplified formats fail to address the complexity of long-horizon tasks that require complicated actions beyond mere QA.
3 Methodology
To conduct our study, we construct a new hallucination probing set by systematically modifying an existing embodied AI benchmarkĀ Li etĀ al. (2024b) to introduce scene-task inconsistencies.
3.1 Preliminaries
Embodied Settings. Let denote the user-provided natural language task description and the corresponding visual observation captured by the agentās camera about the environment. Let denote the set of objects present in the scene and the universal set of all states that any object can attain. A perception module transforms the visual input into a structured linguistic representation that describes the scene in terms of tuples of the form , where is the -th object, is its initial state, and is the set of all possible states the object may attain. This representation captures both the current scene information and the potential state transitions available for objects, as formalized in EquationĀ 1.
An LLM-based goal interpretation module , parameterized by , takes as input the combined context and outputs a structured goal specification in Linear Temporal Logic (LTL)Ā Li etĀ al. (2024b), consisting of a set of node goals and edge goals, as illustrated in FigureĀ 2 and EquationĀ 2. A node goal specifies that object should attain the goal state . An edge goal is represented as a triplet , indicating that object is expected to hold the semantic relationship with object , where and denotes relation (e.g., on, inside, next to).
These symbolic goals are then passed to downstream modules such as action planning and trajectory generationĀ Li etĀ al. (2024b). We highlight that any misinterpretation of the LTL goal is critical, as errors at this stage will propagate through the downstream modules and impair task execution.
Hallucination. Hallucination is commonly defined as an apparent perception in the absence of an external stimulusĀ Ji etĀ al. (2023). In the context of LLM-conditioned embodied agents, we define hallucination as the generation of content that is not grounded in the observed environment Ā Rawte etĀ al. (2023b); Huang etĀ al. (2025), i.e., outputs that reflect objects or states inconsistent with the given scene. Let and denote the sets of objects and states mentioned in the modelās output . A hallucination occurs if any mentioned object or state is not supported by the input scene representation āspecifically, as formalized in EquationĀ 3, if any mentioned object does not exist in the scene () or the predicted state is not from the allowed states ().
3.2 Hallucination Instances
Previous studiesĀ Liu etĀ al. (2023); Guan etĀ al. (2024) show that hallucinations arise when the user query references objects absent from the image, suggesting that inconsistencies between the scene and the user-provided instruction are a key trigger. Inspired by this, we extend an existing benchmarkĀ Li etĀ al. (2024b) to systematically understand hallucination patterns by applying targeted transformations to two core components of the base prompt: task description and scene information , to introduce scene-task inconsistencies.
A. Task Description Modification. In this setting, we modify the task description of the base prompt while keeping the scene unchanged, allowing us to examine how variations in user-provided instructions can trigger hallucinations.
[…]
B. Scene Information Modification. We keep the task description fixed and modify the scene information to investigate hallucinations induced by changes in the environment.
[…]
Ground-Truth Preservation. We also aim to understand how hallucinations may affect the success of tasks. Thus, we induce hallucinations within the existing benchmark while ensuring that evaluation remains aligned with the original LTL plans. Specifically, our modified prompts fall into two categories: (i) those maintaining the validity of the original plansāas DistInj and SynonymSubābecause distractors or synonyms should not change the ultimate goal, and (ii) those creating situations where the correct response is to generate no planāsuch as TaskObjRem and SceneTaskConābecause required objects are missing. This design allows us to reuse the original ground truth for direct comparison and maintain consistent evaluation criteria.
Env | Models | Base Prompt | TaskDescMod | SceneMod | |||||||||||
DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||||
V I R T U A L H O M E | Llama-3-8B-Instruct | 3.7 | 1.2 | 25.6 | 4.8 | 60.51 | 20.4 | 3.7 | 36.60 | 17.0 | 2.6 | 27.36 | 38.5 | 8.7 | 66.45 |
DS-R1-Distil-LLaMA-8B | 19.7 | 4.7 | 48.6 | 7.5 | 56.85 | 37.6 | 5.7 | 38.49 | 40.2 | 5.9 | 30.78 | 57.1 | 9.3 | 63.36 | |
Gemma-2-9b-it | 4.4 | 5.4 | 27.1 | 9.9 | 81.02 | 31.7 | 8.7 | 46.56 | 16.2 | 6.1 | 28.34 | 52.6 | 15.0 | 66.78 | |
Qwen-14B-Instruct | 3.6 | 4.2 | 26.4 | 4.4 | 72.69 | 36.6 | 6.0 | 47.59 | 13.1 | 4.8 | 17.10 | 50.6 | 8.0 | 70.68 | |
DS-R1-Distil-Qwen-14B | 5.6 | 2.9 | 28.2 | 3.7 | 50.76 | 33.4 | 5.0 | 39.86 | 14.6 | 3.9 | 26.22 | 56.5 | 7.3 | 69.87 | |
Llama-4-Scout-17B-16E-Instruct | 1.6 | 0.7 | 6.3 | 3.3 | 18.88 | 21.1 | 9.2 | 36.25 | 14.3 | 1.3 | 26.38 | 48.2 | 8.9 | 67.26 | |
Llama-3.3-70B-Instruct | 1.5 | 0.2 | 12.8 | 0.2 | 21.02 | 22.8 | 1.9 | 46.74 | 8.0 | 1.0 | 21.82 | 37.6 | 0.8 | 60.10 | |
Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 8.27 | 3.0 | 0.0 | 7.21 | 6.0 | 0.0 | 20.08 | |
Gemini 2.5 Flash | 0.5 | 0.0 | 0.6 | 0.0 | 0.2 | 22.4 | 0.0 | 50.86 | 5.7 | 0.0 | 7.82 | 36.0 | 0.0 | 61.89 | |
Gemini 2.5 Flash (w/o thinking) | 1.3 | 0.0 | 2.1 | 1.0 | 5.18 | 14.4 | 0.8 | 39.00 | 12.7 | 0.0 | 20.03 | 18.4 | 0.0 | 47.07 | |
Mistral-Large-Instruct-2411 | 1.6 | 1.0 | 14.0 | 1.4 | 8.73 | 21.1 | 1.5 | 54.64 | 15.2 | 1.6 | 17.10 | 47.1 | 2.1 | 67.75 | |
Claude 3.5 Sonnet | 2.1 | 0.1 | 2.2 | 0.3 | 0.2 | 16.6 | 0.2 | 45.53 | 31.0 | 0.0 | 36.64 | 14.4 | 1.0 | 40.07 | |
GPT-4o | 1.0 | 1.2 | 2.9 | 2.0 | 3.45 | 7.6 | 1.8 | 31.10 | 22.6 | 1.3 | 27.04 | 36.8 | 1.4 | 60.91 | |
B E H A V I O R | Llama-3-8B-Instruct | 1.93 | 2.15 | 8.2 | 10.8 | 6.40 | 5.0 | 6.7 | 14.34 | 2.2 | 7.4 | 0.70 | 17.8 | 7.7 | 18.27 |
DS-R1-Distil-LLaMA-8B | 1.26 | 8.16 | 13.9 | 27.2 | 25.25 | 6.8 | 13.3 | 28.84 | 20.4 | 10.1 | 21.39 | 60 | 19.6 | 25.65 | |
Gemma-2-9b-it | 0.87 | 1.8 | 11.6 | 5.7 | 9.43 | 10.8 | 4.8 | 22.80 | 11.4 | 2.8 | 5.74 | 73.1 | 8.2 | 19.74 | |
Qwen-14B-Instruct | 0.41 | 3.33 | 15.8 | 5.7 | 13.13 | 6.6 | 4.6 | 18.48 | 1.7 | 4.1 | 1.57 | 63.1 | 5.7 | 20.76 | |
DS-R1-Distil-Qwen-14B | 0.32 | 1.24 | 8.4 | 5.2 | 13.13 | 4.5 | 6.0 | 19.17 | 2.8 | 7.6 | 4.52 | 46.4 | 4.7 | 17.99 | |
Llama-4-Scout-17B-16E-Instruct | 0.0 | 1.7 | 0.3 | 2.3 | 0.34 | 2.6 | 1.7 | 11.40 | 3.9 | 1.0 | 3.48 | 32.0 | 1.7 | 15.22 | |
Llama-3.3-70B-Instruct | 0.0 | 2.1 | 1.9 | 2.8 | 2.69 | 3.2 | 1.7 | 15.89 | 0.0 | 1.8 | 0.0 | 45.0 | 1.3 | 12.82 | |
Gemini 2.0 Flash | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 1.1 | 0.2 | 4.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.6 | 4.84 | 0.0 | 0.0 | 0.0 | 12.6 | 1.1 | 5.54 | |
Gemini 2.5 Flash (w/o thinking) | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.4 | 1.4 | 6.22 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | |
Mistral-Large-Instruct-2411 | 0.0 | 0.0 | 0.2 | 1.6 | 0.34 | 2.8 | 0.4 | 12.95 | 0.0 | 0.5 | 0.0 | 38.6 | 0.0 | 12.55 | |
Claude 3.5 Sonnet | 0.0 | 2.0 | 0.0 | 3.0 | 0.0 | 1.0 | 1.2 | 3.28 | 0.0 | 2.5 | 0.0 | 0.0 | 2.8 | 0.0 | |
GPT-4o | 0.0 | 0.0 | 1.9 | 0.5 | 2.36 | 2.6 | 0.0 | 10.71 | 0.0 | 0.0 | 0.0 | 31.9 | 0.0 | 9.87 |
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate long-horizon tasks that require multiple sequential steps across two simulation environments: VirtualHomeĀ Puig etĀ al. (2018) and BEHAVIORĀ Li etĀ al. (2023a). TableĀ 11 in the Appendix shows the prompt distribution in our probing set, which contains 2,574 samples designed to demonstrate task-scene inconsistencies.
Models. We evaluate 12 open and closed LLMs, including models from LLaMAĀ Grattafiori etĀ al. (2024), QwenĀ Yang etĀ al. (2024a), GemmaĀ Team etĀ al. (2024), GeminiĀ Team etĀ al. (2023), ClaudeĀ Anthropic (2024), MistralĀ AI (2024), GPT-4oĀ Achiam etĀ al. (2023), as well as DeepSeek-R1Ā Guo etĀ al. (2025) distilled versions of LLaMA and Qwen. To assess the role of vision, we evaluate VLMs with the same language backbones, including Qwen-VLĀ Wang etĀ al. (2024a), Gemma-3Ā Team etĀ al. (2025), and LLaMA-Vision. Smaller models are more suitable for running locally in robots, while large models represent the state-of-the-art. See TableĀ 12 for model cards.
Metrics. We adopt two widely used metrics originally proposed for image captioning: CHAIR (Caption Hallucination Assessment with Image Relevance)Ā Rohrbach etĀ al. (2018) and POPE (Polling-based Object Probing Evaluation)Ā Li etĀ al. (2023b). Although used for visual grounding, both generalize naturally to text generation. CHAIR, defined in EquationĀ 8, quantifies the proportion of hallucinated items relative to all mentioned ones.
While CHAIR gives a holistic estimate, it can be biased by output length. For instance, if two outputs hallucinate the same number of entities but differ in length, CHAIR may unfairly penalize the shorter one. Therefore, we also report POPE, which frames hallucination detection as binary classification.
Here, we measure whether non-existent objects appear in generated text by posing yes/no questions (e.g., āIs a washing machine mentioned?ā). Since we inject controlled hallucinations, we create a set of binary questions on those non-existent objects. POPE, formalized in EquationĀ 9, computes the proportion of these questions that incorrectly receive a āyesā response. However, for object states and base prompts, we do not have a predefined list of non-existent elements. Therefore, we report POPE for objects and omit it for states and base prompts.
4.2 Results
Models and Environments. Overall, hallucination rates are lower in BEHAVIOR compared to VirtualHome. Among all models, Claude and Gemini are most resistant to hallucination, while smaller models hallucinate more frequently.
Image | Text | Image + Text | |
Qwen-VL-7B-Instruct | 31.84 | 30.32 | 24.37 |
LLaMA-11B-Vision-Instruct | 46.90 | 41.49 | 36.44 |
Gemma-3-12b-it | 36.67 | 32.99 | 24.14 |
Models | Mitigation Method | TaskDescMod | SceneMod | ||||||||||
DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||
Llama-3-8B-Instruct | KAF | 9.0 | 1.4 | 45.99 | 10.9 | 2.8 | 30.93 | 7.9 | 1.6 | 24.76 | 19.4 | 3.3 | 59.12 |
SCW | 2.3 | 1.5 | 0.0 | 7.9 | 2.6 | 5.15 | 4.8 | 1.8 | 7.33 | 17.2 | 5.4 | 25.41 | |
Gemma-2-9b-it | KAF | 10.3 | 5.3 | 65.48 | 17.6 | 8.5 | 42.96 | 9.9 | 4.1 | 25.08 | 38.8 | 9.9 | 60.59 |
SCW | 4.4 | 6.5 | 5.28 | 11.5 | 8.6 | 19.59 | 6.7 | 5.3 | 12.05 | 38.2 | 10.9 | 34.69 | |
Qwen-14B-Instruct | KAF | 12.1 | 2.8 | 47.72 | 21.4 | 5.0 | 41.24 | 5.3 | 2.2 | 10.10 | 18.9 | 4.9 | 55.37 |
SCW | 2.0 | 2.6 | 0.30 | 18.7 | 6 | 17.35 | 5.0 | 3.3 | 3.75 | 20.8 | 4.2 | 42.35 |
Sceneātask Inconsistency Types. Among the four settings, the SceneTaskCon yields the highest hallucination, indicating that models generate plans even in irrelevant scenarios where the ideal response would be to abstain from planning Ā Zhang etĀ al. (2024a); Liu etĀ al. (2023). In contrast, the SynonymSub setting results in the overall lowest hallucination, suggesting that models are generally capable of reasoning synonymous references. Hallucination under the TaskObjRemo falls in the middle ā typically arising when core objects are removed, while the absence of peripheral items has less effect. As illustrated in FigureĀ 2, when both the soap and washing machine are removed from the scene, models tend to hallucinate the washing machine but not the soap, indicating a strong co-occurrence biasĀ Ji etĀ al. (2023); Huang etĀ al. (2025). For DistInj, smaller models are more prone to hallucinating distractors, whereas larger models often ignore them and generate grounded plans.
State Hallucination. State hallucinations stay relatively low, consistent with the fact that we do not explicitly introduce state inconsistencies. A moderate increase in state hallucination compared to the base prompts indicates secondary effects: object hallucinations may also trigger incorrect state predictions. See AppendixĀ F for examples.
Cross-Modality. We analyze the impact of cross-modal scene information on hallucinations (details in AppendixĀ D). We observe that VLMs achieve better grounding when scene information is given in text form compared to image-only, underscoring the visual grounding limitationsĀ Rahmanzadehgervi etĀ al. (2024); Wang etĀ al. (2023b, a). Combining image and text further reduces hallucination by cross-modal verification (see TableĀ 2).
Mitigation. We implement post-hoc self-correctionĀ Madaan etĀ al. (2023); Wang etĀ al. (2024b) as mitigation by prompting models to revise initial responses after receiving feedback. We explore two approaches: (i) Knowledge-Augmented Feedback (KAF)Ā Peng etĀ al. (2023), which provides general guidance, and (ii) Self-Correcting Woodpecker (SCW)Ā Yin etĀ al. (2024), which offers explicit feedback by naming hallucinated objects. As shown in TableĀ 3, SCW outperforms KAF by providing more actionable feedback. However, hallucination rates in the SceneTaskCon setting remain high, highlighting the need for stronger mitigation strategiesĀ Tian etĀ al. (2023); Elaraby etĀ al. (2023). Further details and prompt formats are provided in AppendixĀ C.
5 Discussion
Type | Task Description | Scene Objects | Predicted LTL Goal |
DistInj |
Place books on the shelf. Be sure not to bump into the printer.
# Injected printer as distractor |
book.n.02_1, book.n.02_2,
shelf.n.01_1 |
{ā¦"not", "Touching", "agent.n.01_1", "printer.n.01_1"ā¦} |
TaskObjRem | Place the books into the carton for storage |
floor.n.01_1, shelf.n.01_1, carton.n.02_1
# Romoved book from the scene |
{ā¦["inside", "book.n.01_1", "carton.n.02_1"], ["inside", "book.n.01_2", "carton.n.02_1"]ā¦} |
SynonymSub | Leave the soap in the bucket and place the bucket next to the sink |
detergent.n.01_1, pail.n.01_1, basin.n.01_1
# Replaced scene objects with synonyms |
{ ā¦[["NextTo", "bucket.n.01_1", "basin.n.01_1"], ["Inside", "detergent.n.01_1", "pail.n.01_1"]..} |
SceneTaskCon | Place book on the shelf |
scanner.n.02_1, table.n.02_1, floor.n.01_1
# Replaced with contradictory scene |
{ā¦["ontop", "book.n.01_1","shelf.n.01_1"]]ā¦} |
5.1 Understanding Hallucination
We identify three key contributing factors under which hallucinations emerge: task complexity, hallucination variants, and model capability.
Task Complexity. Tasks in VirtualHome are significantly more complex than those in BEHAVIOR, often requiring abstract scene understandingāfactors that contribute to overall higher hallucination in VirtualHome (see TableĀ 1). As shown in AppendixĀ FigureĀ 3, BEHAVIOR task descriptions tend to be low-level and closely aligned with symbolic LTL goals, typically involving direct pick-and-place actions (e.g., placing a modem under a table). In contrast, VirtualHome task descriptionsāsuch as writing an emailārequire understanding of the scene, involving turning on a computer, holding the keyboard, and so on. These gaps between the task description and observable scene lead the model to infer and āfill inā missing steps or objects. Although this reflects a form of reasoning, it often leads to hallucinations, highlighting that LLMs still struggle to reliably execute long-horizon tasks.
Hallucination Variants. As shown in TableĀ 4, hallucinations arise from the modelās attempt to complete the task despite missing inputs. Models do not simply copy the task description; instead, there is evidence of underlying reasoning and an attempt to resolve inconsistencies, often by filling in gaps based on learned patterns. In the SynonymSub, we observe that models recognize that substituted objects conceptually belong to the same category (e.g., āpailā and ābucketā are the same), yet they fail to maintain naming consistency in the output. This inconsistency may stem from language modeling biases: in conversational settings, synonym variation is preferred for fluency, but in task-oriented planningāespecially grounded in a specific sceneāconsistent naming is critical for accurate symbolic goal generation. In particular, these patterns suggest that while LLMs exhibit signs of reasoning, they lack the control mechanisms necessary to balance user instructions with the observed scene.
Model Capability.
Task Description | Scene Objects | Predicted LTL Goal |
Turn on light | table, cupboard, plate | I cannot fulfill this request. Thereās not enough information about light. |
Pick up phone |
stereo,
trashcan |
Given the goal "Pick up phone," the phone is not in the provided object list. Therefore, this goal cannot be translated to a symbolic goal. āājson{"node goals": [],"edge goals": [],}āā |
Smaller models hallucinate more than larger onesāsuggesting that increased scale causes reduced hallucination. As shown in TableĀ 5, Gemini-2.0 Flash is the only model among our studied ones to reject generation when required objects are missing; others generate empty plans. Seeing the stronger performance of larger models, we evaluate their behavior in non-hallucinated cases (see AppendixĀ A). We find that plan quality remains stable under DistInj and SynonymSub. However, under TaskObjRem and SceneTaskCon, these models are less likely to generate empty plansāhighlighting that even larger models struggle in handling task infeasibility. We discuss the challenge of grounding reasoning in AppendixĀ B.
5.2 Beyond Hallucinations: Planning Failures
Task Description | Scene Objects | Predicted LTL Goal |
ā¦Turn on washing machine. | bathroom, shower | {ā¦{"name": "shower","state": "ON"}]⦠|
Pick up cat.
Rub hand on cat. |
washing_machine, clothes_pants, | There is no cat in the environment, so this goal cannot be fulfilled. I will set the goal to pick up clothes and hold it. ānode goalsā: [{ānameā: āclothes_pantsā, āstateā: āGRABBEDā] |
Wax the dust off the vehicle. |
bed.n.01_1,
door.n.01_1 |
ā¦"not",["Dusty",
"bed.n.01_1"]], ["not", ["Dusty","door.n.01_1"] |
While hallucination is a failure mode, we observe that its absence does not guarantee correct planning. Even when models do not hallucinate, they often generate syntactically valid but incorrect LTL goals (see TableĀ 6). These models attempt to fulfill the task by re-purposing available scene objects inappropriately. It highlights a deeper issue that the models still struggle to recognize when not to generate plans. See AppendixĀ E on reliable planning.
6 Conclusion
In this paper, we systematically identify hallucination cases in LLM-based embodied agents, offering insights into where and how current models fall short when executing long-horizon tasks under sceneātask inconsistencies. The significantly elevated hallucination rates observed with our modified prompts highlight that existing models struggle to reconcile mismatches between user instructions and the environment. We further find that hallucinations persist despite feedback-based mitigation and are only partially reduced with cross-modal inputs. By introducing challenging scenarios along with guidance on ideal model behavior, our work takes an important first step toward enabling more grounded planning in real-world embodied agents.
7 Limitations
While our empirical study reveals significant occurrences of hallucinations in LLM-based embodied agents under sceneātask inconsistencies, it has some limitations. First, although our prompt modifications introduce diverse and challenging failure cases, they do not exhaustively cover the full space of hallucination types in embodied agents. Our focus is primarily on object-level hallucinations, but other formsāsuch as hallucinations related to counts, attributes, or spatial relationsāremain unexplored. Second, our analysis is limited to symbolic outputs (in the form of Linear Temporal Logic plans) and does not evaluate downstream execution errors that may arise during physical task performance in real-world environments. Third, our cross-modal verification experiments are conducted at a small scale and limited to the Distractor Injection. Extending such experiments to other modified prompts would require changing the underlying simulation environments to attain the scene information as images, which we leave for future work. Fourth, we do not evaluate all the available models, so our sampling may not be representative enough for the broader landscape of LLM capabilities. Lastly, we explore only post-hoc, self-correction-based mitigation strategies. Other mitigation approachesāsuch as alternative decoding methods, supervised fine-tuning, or architecture-level modificationsāmay yield deeper insights and are left for future exploration.
Appendix A Goal Interpretation Correctness in Non-Hallucinated Cases
We evaluate model performance in non-hallucinated cases. For DistInj and SynonymSub, the ideal output is the original LTL goal. Using the original benchmarkās evaluation script, we report performance in TableĀ 7, showing that larger models maintain performance in these settings. For TaskObjRem and SceneTaskCon, the ideal response is to reject the task or generate an empty plan. We report refusal/empty response rates in TableĀ 8, revealing that models still struggle to recognize and appropriately handle task infeasibility.
Appendix B Reasoning Models
While reasoning-capable models are expected to be less prone to hallucination, our results (TableĀ 1) show mixed performance for smaller modelsāsometimes outperforming the non-reasoning version of model, and sometimes underperforming. To better understand this behavior, we examine the reasoning traces (i.e., āthinkingā blocks) of smaller models such as DeepSeek-R1 distilled versions of LLaMA and Qwen. As shown in TableĀ 9, we find that these models often attempt to fill in missing information by assuming the existence of objects not present in the scene. For example, when instructed to interact with an object like a āwashing machineā or āTVā that is not listed in the scene, they rationalize its presence based on task relevance and proceed to generate symbolic goals involving it. While such reasoning demonstrates an ability to infer implicit dependencies, it ultimately leads to incorrect grounding. Enabling reliable reasoning for grounded planningāespecially in small modelsāremains an open problem.
Task Description | Scene Objects | Models | Reasoning |
Open washing machine door. Put dirty clothes into washing machine. Measure laundry detergent and place into washing machine. Turn dial to appropriate wash setting. Power on. | table, cupboard, plate | DS-R1-Distil-LLaMA-8B | <think> ⦠So āwashing machineā isnāt listed, which is a problem. Wait, perhaps I need to include it as an object. But in the initial objects, itās not there. Maybe itās implied but not listed, so I should proceed without it or assume itās part of the sceneā¦So, perhaps the washing machine is part of the scene ...</think> |
DS-R1-Distil-Qwen-14B | <think>⦠Wait, actually, the washing machine isnāt listed in the initial objects, but itās implied because the task involves loading clothes into it. So I should probably include it as a relevant object.ā¦</think> | ||
I find TV remote. I grab TV remote. I sit on sofa. I switch on TV. I switch channel. |
coffee_maker,
coffee_filter, ground_coffee |
DS-R1-Distil-LLaMA-8B | <think> ⦠Wait, but the TV isnāt listed as a relevant object, so I canāt include it in the node goals. Thatās a problem. Maybe I should proceed by including the TV as a node goal if possible, even if itās not in the initial data. Iām not sure. I think Iāll proceed by including it as a node goal with its state as ON, assuming itās part of the scene, even if itās not explicitly listed.ā¦</think> |
DS-R1-Distil-Qwen-14B | <think> ⦠The TVās initial state isnāt specified, but itās probably on since the user wants to change the channel. The remoteās initial state is also unclear, but itās likely either clean or in a state where it can be used, maybe plugged in or just lying around. ā¦</think> | ||
Turn on light, grab a book, start to read the book | computer, chair, keyboard | DS-R1-Distil-LLaMA-8B | <think> ⦠I need to find the correct state for the light. Looking at the possible states for ālightā isnāt provided, but in general, turning it on would mean itās āONā. So, a node goal would be the light being āONāā¦.</think> |
DS-R1-Distil-Qwen-14B | <think> ā¦Assuming the light is an object, I need to set its state to āONā.But since itās not listed, maybe I need to add it.ā¦</think> |
Appendix C Mitigation Details
Knowledge-Augmented Feedback (KAF) In KAFĀ Peng etĀ al. (2023), we provide general feedback by appending the prior response and instructing the model with: "Your previous response is inconsistent with the scene information. Please refine your answer to the prior request based on your examination."
Self-Correcting Woodpecker (SCW). In SCWĀ Yin etĀ al. (2024), we provide targeted feedback by listing hallucinated objects explicitly: "The scene does not contain {objects}. Please refine your answer to the prior request based on your examination." This approach makes the feedback more specific and actionable, leading to greater hallucination reduction.
Appendix D Cross Modal Experiment Details
We conduct a small-scale analysis of cross-modal hallucination behavior using the distractor injection setting in VirtualHome. In this setting, only the task description is modified, while the scene remains unchanged. Therefore, we are able to directly use the original simulator-provided scene images for evaluation. Extending cross-modal analysis to other variants in our hallucination probing set would require modifying the visual scenes within the simulatorāan effort we leave for future work.
Appendix E Toward Reliable Planning
We also provide guidance on what should or should not be done for reliable planning in each scenario of our hallucination probing set. In embodied settings, hallucinating an object that is not present in the scene can cause the agent to search for or interact with nonexistent entitiesāleading to navigation errors, execution failures, or unsafe behaviors. With our setup, model behavior should ideally fall into one of two expected response patterns: (i) when the task is feasible and the required objects are present (as in DistInj and SynonymSub), the model should generate a plan that aligns with the original ground-truth LTL goal; (ii) when key objects are missing or the task is semantically incompatible with the scene (as in TaskObjRem and SceneTaskCon), the model should abstain from generating a plan altogether.
Appendix F State Hallucinations
In addition to object-level hallucinations, we observe cases where models generate invalid or unsupported object states in an attempt to satisfy task instructions. These occur when the predicted state does not belong to the predefined set of possible states for the given object. As shown in TableĀ 10, models sometimes fabricate statesāsuch as āARRANGEDā or āINSIDEāāthat are not included in the allowed state space.
Task Description | Scene Objects | Predicted LTL Goal |
Arrange chairs, set napkin, set plate, set knife, set fork and glass on table. | ā¦chair, initial states: [āCLEANā], possible states: [āCLEANā, āFREEā, āGRABBEDā, āOCCUPIEDā]⦠|
⦠{""name"": ""chair"",
""state"": ""ARRANGED""}⦠|
Find coffee maker, find coffee filter and place it in coffee maker. Find ground coffee and water. put both in coffee maker. Close the coffee maker and switch it on. | ā¦coffee_filter, initial states: [āCLEANā], possible states: [āCLEANā, āDIRTYā, āCLEANā, āDIRTYā, āGRABBEDā]⦠|
ā¦{""name"": ""coffee_filter"",
""state"": ""INSIDE coffee_maker""}⦠|
I will load the dirty clothes into the washing machine. | ā¦clothes_pants, initial states: [āCLEANā], possible states: [āCLEANā, āDIRTYā, āCLEANā, āDIRTYā, āFOLDEDā, āFREEā, āGRABBEDā, āOCCUPIEDā, āUNFOLDEDā]⦠|
⦠{""name"": ""clothes_pants"",
""state"": ""INSIDE""}⦠|
Appendix G Hallucination Probing Prompts Design and Distribution
G.1 Base Prompt Format and Structure
The base prompts from VirtualHome are more complex, often requiring multi-step reasoning about scene objects and their interactions. In contrast, BEHAVIOR prompts closely resemble their corresponding LTL goals, leaving less room for ambiguity and reducing the modelās need to infer or fill in missing information.
G.2 Prompt Distribution
We report the total number of prompts generated for each hallucination variantāDistractor Injection (DistInj), Task-Relevant Object Removal (TaskObjRem), Synonymous Object Substitution (SynonymSub), and Scene Task Contradiction (SceneTaskCon)āwithin both VirtualHome and BEHAVIOR environments. In total, our hallucination probing set contains 2,574 examples designed to evaluate model robustness under task-scene inconsistencies.
Env | Transformation Type | |N| |
VirtualHome | DistInj | 338 |
TaskObjRem | 582 | |
SynonymSub | 338 | |
SceneTaskCon | 338 | |
Behavior | DistInj | 100 |
TaskObjRem | 678 | |
SynonymSub | 100 | |
SceneTaskCon | 100 | |
Total | 2574 |
G.3 GPT-4o Prompts
FigureĀ 4 and FigureĀ 5 show the prompts used to query GPT-4o to generate our halluciantion probing set.
Appendix H Model Cards
TableĀ 12 shows the models cards used in our experiments.
Model Name | Complete Model ID | Hosting |
Llama-3-8B-Instruct | meta-llama/Meta-Llama-3-8B-Instruct | Hugging Face |
DS-R1-Distil-LLaMA-8B | deepseek-ai/DeepSeek-R1-Distill-Llama-8B | Hugging Face |
Gemma-2-9b-it | google/gemma-2-9b-it | Hugging Face |
Qwen-14B-Instruct | Qwen/Qwen2.5-14B-Instruct | Hugging Face |
DS-R1-Distil-Qwen-14B | deepseek-ai/DeepSeek-R1-Distill-Qwen-14B | Hugging Face |
Llama-4-Scout-17B-16E-Instruct | llama-4-scout-17b-16e-instruct-maas | GCP Vertex |
Llama-3.3-70B-Instruct | llama-3.3-70b-instruct-maas | GCP Vertex |
Gemini 2.0 Flash | gemini-2.0-flash-001 | GCP Vertex |
Gemini 2.5 Flash | gemini-2.5-flash-preview-04-17 | GCP Vertex |
Mistral-Large-Instruct-2411 | mistral-large-instruct-2411 | GCP Vertex |
Claude 3.5 Sonnet | claude-3-5-sonnet-v2@20241022 | GCP Vertex |
Qwen-VL-7B-Instruct | Qwen/Qwen2.5-VL-7B-Instruct | Hugging Face |
LLaMA-11B-Vision-Instruct | meta-llama/Llama-3.2-11B-Vision-Instruct | Hugging Face |
Gemma-3-12b-it | google/gemma-3-12b-it | Hugging Face |

### v2

@@
 mbodied gents Driven by arge Language Models
-Trishna ChakrabortyāUdita Ghosh āXiaopan ZhangāFahim Faisal Niloy Yue Dong āJiachen Li āAmit K.
-Roy-Chowdhury āChengyu Song University of California, Riverside {tchak006,ughos002,xzhan006,fnilo001,jiachen.li,yue.dong}@ucr.edu csong@cs.ucr.edu āamitrc@ece.ucr.edu
+Trishna ChakrabortyāāUdita Ghosh āāXiaopan ZhangāāFahim Faisal Niloy Yue Dong āJiachen Li āAmit K.
+Roy-Chowdhury āChengyu Song University of California, Riverside {tchak006, ughos002, xzhan006, fnilo001, yue.dong, jiachen.li}@ucr.edu āamitrc@ece.ucr.edu ācsong@cs.ucr.edu
 1 Introduction
@@
 To fill this gap, we present, to the best of our knowledge, the first empirical study that systematically exemplifies and quantifies hallucinations in long-horizon planning through controlled scene-task inconsistencies.
-Because existing datasets either do not explicitly aim to elicit hallucination in long-horizon embodied tasks, or do not target embodied agent tasks, we first construct a new probing set that is more likely to cause hallucination.
+Because existing datasets either do not explicitly aim to elicit hallucination in long-horizon embodied tasks, or do not target embodied agent tasks, we first construct a new probing set111The HEAL probing set is publicly available at https://huggingface.co/datasets/Trishna13/HEAL that is more likely to cause hallucination.
 Constructing such a probing set faces two main challenges: (i) how to induce hallucinations and (ii) how to detect unwanted behaviors caused by hallucinations.
@@
 Hallucinations in LLMs.
-Hallucinations in LLMs have been widely studied in text-onlyĀ Dhuliawala etĀ al.
+Hallucinations in LLMs have been widely studied in text-only settingsĀ Dhuliawala etĀ al.
 (2023); Agrawal etĀ al.
 (2024).
-Recent works extend these to multimodal LLMs Ā Bai etĀ al.
+Recent work extends these to multimodal LLMs, manifested as category misidentification, attribute errors, or relation misrepresentation Ā Bai etĀ al.
 (2024); Xu etĀ al.
@@
 This design allows us to reuse the original ground truth for direct comparison and maintain consistent evaluation criteria.
-Env | Models | Base Prompt | TaskDescMod | SceneMod | |||||||||||
-DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||||
-V I R T U A L H O M E | Llama-3-8B-Instruct | 3.7 | 1.2 | 25.6 | 4.8 | 60.51 | 20.4 | 3.7 | 36.60 | 17.0 | 2.6 | 27.36 | 38.5 | 8.7 | 66.45 |
-DS-R1-Distil-LLaMA-8B | 19.7 | 4.7 | 48.6 | 7.5 | 56.85 | 37.6 | 5.7 | 38.49 | 40.2 | 5.9 | 30.78 | 57.1 | 9.3 | 63.36 | |
-Gemma-2-9b-it | 4.4 | 5.4 | 27.1 | 9.9 | 81.02 | 31.7 | 8.7 | 46.56 | 16.2 | 6.1 | 28.34 | 52.6 | 15.0 | 66.78 | |
-Qwen-14B-Instruct | 3.6 | 4.2 | 26.4 | 4.4 | 72.69 | 36.6 | 6.0 | 47.59 | 13.1 | 4.8 | 17.10 | 50.6 | 8.0 | 70.68 | |
-DS-R1-Distil-Qwen-14B | 5.6 | 2.9 | 28.2 | 3.7 | 50.76 | 33.4 | 5.0 | 39.86 | 14.6 | 3.9 | 26.22 | 56.5 | 7.3 | 69.87 | |
-Llama-4-Scout-17B-16E-Instruct | 1.6 | 0.7 | 6.3 | 3.3 | 18.88 | 21.1 | 9.2 | 36.25 | 14.3 | 1.3 | 26.38 | 48.2 | 8.9 | 67.26 | |
-Llama-3.3-70B-Instruct | 1.5 | 0.2 | 12.8 | 0.2 | 21.02 | 22.8 | 1.9 | 46.74 | 8.0 | 1.0 | 21.82 | 37.6 | 0.8 | 60.10 | |
-Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 8.27 | 3.0 | 0.0 | 7.21 | 6.0 | 0.0 | 20.08 | |
-Gemini 2.5 Flash | 0.5 | 0.0 | 0.6 | 0.0 | 0.2 | 22.4 | 0.0 | 50.86 | 5.7 | 0.0 | 7.82 | 36.0 | 0.0 | 61.89 | |
-Gemini 2.5 Flash (w/o thinking) | 1.3 | 0.0 | 2.1 | 1.0 | 5.18 | 14.4 | 0.8 | 39.00 | 12.7 | 0.0 | 20.03 | 18.4 | 0.0 | 47.07 | |
-Mistral-Large-Instruct-2411 | 1.6 | 1.0 | 14.0 | 1.4 | 8.73 | 21.1 | 1.5 | 54.64 | 15.2 | 1.6 | 17.10 | 47.1 | 2.1 | 67.75 | |
-Claude 3.5 Sonnet | 2.1 | 0.1 | 2.2 | 0.3 | 0.2 | 16.6 | 0.2 | 45.53 | 31.0 | 0.0 | 36.64 | 14.4 | 1.0 | 40.07 | |
-GPT-4o | 1.0 | 1.2 | 2.9 | 2.0 | 3.45 | 7.6 | 1.8 | 31.10 | 22.6 | 1.3 | 27.04 | 36.8 | 1.4 | 60.91 | |
-B E H A V I O R | Llama-3-8B-Instruct | 1.93 | 2.15 | 8.2 | 10.8 | 6.40 | 5.0 | 6.7 | 14.34 | 2.2 | 7.4 | 0.70 | 17.8 | 7.7 | 18.27 |
-DS-R1-Distil-LLaMA-8B | 1.26 | 8.16 | 13.9 | 27.2 | 25.25 | 6.8 | 13.3 | 28.84 | 20.4 | 10.1 | 21.39 | 60 | 19.6 | 25.65 | |
-Gemma-2-9b-it | 0.87 | 1.8 | 11.6 | 5.7 | 9.43 | 10.8 | 4.8 | 22.80 | 11.4 | 2.8 | 5.74 | 73.1 | 8.2 | 19.74 | |
-Qwen-14B-Instruct | 0.41 | 3.33 | 15.8 | 5.7 | 13.13 | 6.6 | 4.6 | 18.48 | 1.7 | 4.1 | 1.57 | 63.1 | 5.7 | 20.76 | |
-DS-R1-Distil-Qwen-14B | 0.32 | 1.24 | 8.4 | 5.2 | 13.13 | 4.5 | 6.0 | 19.17 | 2.8 | 7.6 | 4.52 | 46.4 | 4.7 | 17.99 | |
-Llama-4-Scout-17B-16E-Instruct | 0.0 | 1.7 | 0.3 | 2.3 | 0.34 | 2.6 | 1.7 | 11.40 | 3.9 | 1.0 | 3.48 | 32.0 | 1.7 | 15.22 | |
-Llama-3.3-70B-Instruct | 0.0 | 2.1 | 1.9 | 2.8 | 2.69 | 3.2 | 1.7 | 15.89 | 0.0 | 1.8 | 0.0 | 45.0 | 1.3 | 12.82 | |
-Gemini 2.0 Flash | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 1.1 | 0.2 | 4.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |
-Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.6 | 4.84 | 0.0 | 0.0 | 0.0 | 12.6 | 1.1 | 5.54 | |
-Gemini 2.5 Flash (w/o thinking) | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.4 | 1.4 | 6.22 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | |
-Mistral-Large-Instruct-2411 | 0.0 | 0.0 | 0.2 | 1.6 | 0.34 | 2.8 | 0.4 | 12.95 | 0.0 | 0.5 | 0.0 | 38.6 | 0.0 | 12.55 | |
-Claude 3.5 Sonnet | 0.0 | 2.0 | 0.0 | 3.0 | 0.0 | 1.0 | 1.2 | 3.28 | 0.0 | 2.5 | 0.0 | 0.0 | 2.8 | 0.0 | |
-GPT-4o | 0.0 | 0.0 | 1.9 | 0.5 | 2.36 | 2.6 | 0.0 | 10.71 | 0.0 | 0.0 | 0.0 | 31.9 | 0.0 | 9.87 |
+| Env | Models | Base Prompt | TaskDescMod | SceneMod | ||||||||||||||||||||||
+| DistInj | TaskObjRem | SynonymSub | SceneTaskCon | |||||||||||||||||||||||
+Llama-3-8B-Instruct | 3.7 | 1.2 | 25.6 | 4.8 | 60.51 | 20.4 | 3.7 | 36.60 | 17.0 | 2.6 | 27.36 | 38.5 | 8.7 | 66.45 | |||||||||||
+| DS-R1-Distil-LLaMA-8B | 19.7 | 4.7 | 48.6 | 7.5 | 56.85 | 37.6 | 5.7 | 38.49 | 40.2 | 5.9 | 30.78 | 57.1 | 9.3 | 63.36 | ||||||||||||
+| Gemma-2-9b-it | 4.4 | 5.4 | 27.1 | 9.9 | 81.02 | 31.7 | 8.7 | 46.56 | 16.2 | 6.1 | 28.34 | 52.6 | 15.0 | 66.78 | ||||||||||||
+| Qwen-14B-Instruct | 3.6 | 4.2 | 26.4 | 4.4 | 72.69 | 36.6 | 6.0 | 47.59 | 13.1 | 4.8 | 17.10 | 50.6 | 8.0 | 70.68 | ||||||||||||
+| DS-R1-Distil-Qwen-14B | 5.6 | 2.9 | 28.2 | 3.7 | 50.76 | 33.4 | 5.0 | 39.86 | 14.6 | 3.9 | 26.22 | 56.5 | 7.3 | 69.87 | ||||||||||||
+| Llama-4-Scout-17B-16E-Instruct | 1.6 | 0.7 | 6.3 | 3.3 | 18.88 | 21.1 | 9.2 | 36.25 | 14.3 | 1.3 | 26.38 | 48.2 | 8.9 | 67.26 | ||||||||||||
+| Llama-3.3-70B-Instruct | 1.5 | 0.2 | 12.8 | 0.2 | 21.02 | 22.8 | 1.9 | 46.74 | 8.0 | 1.0 | 21.82 | 37.6 | 0.8 | 60.10 | ||||||||||||
+| Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 2.2 | 0.0 | 8.27 | 3.0 | 0.0 | 7.21 | 6.0 | 0.0 | 20.08 | ||||||||||||
+| Gemini 2.5 Flash | 0.5 | 0.0 | 0.6 | 0.0 | 0.2 | 22.4 | 0.0 | 50.86 | 5.7 | 0.0 | 7.82 | 36.0 | 0.0 | 61.89 | ||||||||||||
+| Gemini 2.5 Flash (w/o thinking) | 1.3 | 0.0 | 2.1 | 1.0 | 5.18 | 14.4 | 0.8 | 39.00 | 12.7 | 0.0 | 20.03 | 18.4 | 0.0 | 47.07 | ||||||||||||
+| Mistral-Large-Instruct-2411 | 1.6 | 1.0 | 14.0 | 1.4 | 8.73 | 21.1 | 1.5 | 54.64 | 15.2 | 1.6 | 17.10 | 47.1 | 2.1 | 67.75 | ||||||||||||
+| Claude 3.5 Sonnet | 2.1 | 0.1 | 2.2 | 0.3 | 0.2 | 16.6 | 0.2 | 45.53 | 31.0 | 0.0 | 36.64 | 14.4 | 1.0 | 40.07 | ||||||||||||
+| GPT-4o | 1.0 | 1.2 | 2.9 | 2.0 | 3.45 | 7.6 | 1.8 | 31.10 | 22.6 | 1.3 | 27.04 | 36.8 | 1.4 | 60.91 | ||||||||||||
+Llama-3-8B-Instruct | 1.93 | 2.15 | 8.2 | 10.8 | 6.40 | 5.0 | 6.7 | 14.34 | 2.2 | 7.4 | 0.70 | 17.8 | 7.7 | 18.27 | |||||||||||
+| DS-R1-Distil-LLaMA-8B | 1.26 | 8.16 | 13.9 | 27.2 | 25.25 | 6.8 | 13.3 | 28.84 | 20.4 | 10.1 | 21.39 | 60 | 19.6 | 25.65 | ||||||||||||
+| Gemma-2-9b-it | 0.87 | 1.8 | 11.6 | 5.7 | 9.43 | 10.8 | 4.8 | 22.80 | 11.4 | 2.8 | 5.74 | 73.1 | 8.2 | 19.74 | ||||||||||||
+| Qwen-14B-Instruct | 0.41 | 3.33 | 15.8 | 5.7 | 13.13 | 6.6 | 4.6 | 18.48 | 1.7 | 4.1 | 1.57 | 63.1 | 5.7 | 20.76 | ||||||||||||
+| DS-R1-Distil-Qwen-14B | 0.32 | 1.24 | 8.4 | 5.2 | 13.13 | 4.5 | 6.0 | 19.17 | 2.8 | 7.6 | 4.52 | 46.4 | 4.7 | 17.99 | ||||||||||||
+| Llama-4-Scout-17B-16E-Instruct | 0.0 | 1.7 | 0.3 | 2.3 | 0.34 | 2.6 | 1.7 | 11.40 | 3.9 | 1.0 | 3.48 | 32.0 | 1.7 | 15.22 | ||||||||||||
+| Llama-3.3-70B-Instruct | 0.0 | 2.1 | 1.9 | 2.8 | 2.69 | 3.2 | 1.7 | 15.89 | 0.0 | 1.8 | 0.0 | 45.0 | 1.3 | 12.82 | ||||||||||||
+| Gemini 2.0 Flash | 0.0 | 0.0 | 0.2 | 0.0 | 0.0 | 1.1 | 0.2 | 4.49 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | ||||||||||||
+| Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.6 | 4.84 | 0.0 | 0.0 | 0.0 | 12.6 | 1.1 | 5.54 | ||||||||||||
+| Gemini 2.5 Flash (w/o thinking) | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 1.4 | 1.4 | 6.22 | 0.0 | 0.0 | 0.0 | 0.0 | 1.1 | 0.0 | ||||||||||||
+| Mistral-Large-Instruct-2411 | 0.0 | 0.0 | 0.2 | 1.6 | 0.34 | 2.8 | 0.4 | 12.95 | 0.0 | 0.5 | 0.0 | 38.6 | 0.0 | 12.55 | ||||||||||||
+| Claude 3.5 Sonnet | 0.0 | 2.0 | 0.0 | 3.0 | 0.0 | 1.0 | 1.2 | 3.28 | 0.0 | 2.5 | 0.0 | 0.0 | 2.8 | 0.0 | ||||||||||||
+| GPT-4o | 0.0 | 0.0 | 1.9 | 0.5 | 2.36 | 2.6 | 0.0 | 10.71 | 0.0 | 0.0 | 0.0 | 31.9 | 0.0 | 9.87 | ||||||||||||
 4 Experiments
@@
 (2023a).
-TableĀ 11 in the Appendix shows the prompt distribution in our probing set, which contains 2,574 samples designed to demonstrate task-scene inconsistencies.
+TableĀ 13 in the Appendix shows the prompt distribution in our probing set, which contains 2,574 samples designed to demonstrate task-scene inconsistencies.
 Models.
@@
 Smaller models are more suitable for running locally in robots, while large models represent the state-of-the-art.
-See TableĀ 12 for model cards.
+All experiments are performed with three independent runs per setting, and results reflect the mean performance.
+See TableĀ 14 for model cards.
 Metrics.
@@
 | Gemma-3-12b-it | 36.67 | 32.99 | 24.14 |
-Models | Mitigation Method | TaskDescMod | SceneMod | ||||||||||
-DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||
-Llama-3-8B-Instruct | KAF | 9.0 | 1.4 | 45.99 | 10.9 | 2.8 | 30.93 | 7.9 | 1.6 | 24.76 | 19.4 | 3.3 | 59.12 |
-SCW | 2.3 | 1.5 | 0.0 | 7.9 | 2.6 | 5.15 | 4.8 | 1.8 | 7.33 | 17.2 | 5.4 | 25.41 | |
-Gemma-2-9b-it | KAF | 10.3 | 5.3 | 65.48 | 17.6 | 8.5 | 42.96 | 9.9 | 4.1 | 25.08 | 38.8 | 9.9 | 60.59 |
-SCW | 4.4 | 6.5 | 5.28 | 11.5 | 8.6 | 19.59 | 6.7 | 5.3 | 12.05 | 38.2 | 10.9 | 34.69 | |
-Qwen-14B-Instruct | KAF | 12.1 | 2.8 | 47.72 | 21.4 | 5.0 | 41.24 | 5.3 | 2.2 | 10.10 | 18.9 | 4.9 | 55.37 |
-SCW | 2.0 | 2.6 | 0.30 | 18.7 | 6 | 17.35 | 5.0 | 3.3 | 3.75 | 20.8 | 4.2 | 42.35 |
+| Models |
+TaskDescMod | SceneMod | ||||||||||||
+| DistInj | TaskObjRem | SynonymSub | SceneTaskCon | ||||||||||||
+| Llama-3-8B-Instruct | KAF | 9.0 | 1.4 | 45.99 | 10.9 | 2.8 | 30.93 | 7.9 | 1.6 | 24.76 | 19.4 | 3.3 | 59.12 | ||
+| SCW | 2.3 | 1.5 | 0.0 | 7.9 | 2.6 | 5.15 | 4.8 | 1.8 | 7.33 | 17.2 | 5.4 | 25.41 | |||
+| Gemma-2-9b-it | KAF | 10.3 | 5.3 | 65.48 | 17.6 | 8.5 | 42.96 | 9.9 | 4.1 | 25.08 | 38.8 | 9.9 | 60.59 | ||
+| SCW | 4.4 | 6.5 | 5.28 | 11.5 | 8.6 | 19.59 | 6.7 | 5.3 | 12.05 | 38.2 | 10.9 | 34.69 | |||
+| Qwen-14B-Instruct | KAF | 12.1 | 2.8 | 47.72 | 21.4 | 5.0 | 41.24 | 5.3 | 2.2 | 10.10 | 18.9 | 4.9 | 55.37 | ||
+| SCW | 2.0 | 2.6 | 0.30 | 18.7 | 6 | 17.35 | 5.0 | 3.3 | 3.75 | 20.8 | 4.2 | 42.35 | |||
 Sceneātask Inconsistency Types.
@@
 A moderate increase in state hallucination compared to the base prompts indicates secondary effects: object hallucinations may also trigger incorrect state predictions.
-See AppendixĀ F for examples.
+See subsectionĀ F.1 for examples.
 Cross-Modality.
@@
 Other mitigation approachesāsuch as alternative decoding methods, supervised fine-tuning, or architecture-level modificationsāmay yield deeper insights and are left for future exploration.
+Acknowledgments
+This research was supported and funded by the NSF grants CMMI-2326309, III-1901379, CNS-2312395, 2431569, UC MRPI and SoCalHub.
+The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.
+Finally, we are grateful to the anonymous reviewers for their constructive feedback.
 Appendix A Goal Interpretation Correctness in Non-Hallucinated Cases
@@
 With our setup, model behavior should ideally fall into one of two expected response patterns: (i) when the task is feasible and the required objects are present (as in DistInj and SynonymSub), the model should generate a plan that aligns with the original ground-truth LTL goal; (ii) when key objects are missing or the task is semantically incompatible with the scene (as in TaskObjRem and SceneTaskCon), the model should abstain from generating a plan altogether.
-Appendix F State Hallucinations
+Appendix F Other Forms of Hallucinations
+F.1 State Hallucinations
 In addition to object-level hallucinations, we observe cases where models generate invalid or unsupported object states in an attempt to satisfy task instructions.
@@
 ""state"": ""INSIDE""}⦠|
-Appendix G Hallucination Probing Prompts Design and Distribution
-G.1 Base Prompt Format and Structure
+F.2 Relation Hallucinations
+Following our definition in subsectionĀ 3.1, relation hallucination occurs when the predicted relations fall outside the allowable set.
+As shown in TableĀ 11, smaller models exhibit higher rates of relation hallucination, consistent with trends observed for object and state hallucinations in the main paper.
+We highlight that the relation hallucinations observed here primarily arise as a secondary effect of the object-level inconsistencies introduced in our study.
+Systematically designing scenarios to explicitly induce attribute- or spatial-level hallucinations remains a valuable but distinct direction, which we leave for future work.
+| Models | Base Prompt | DistInj | TaskObjRem | SynonymSub | SceneTaskCon |
+| Llama-3-8B-Instruct | 17.0 | 22.9 | 18.5 | 17.3 | 19.0 |
+| DS-R1-Distil-LLaMA-8B | 19.6 | 30.2 | 19.8 | 23.4 | 23.4 |
+| Gemma-2-9b-it | 3.8 | 4.8 | 8.0 | 6.5 | 3.9 |
+| Qwen-14B-Instruct | 13.1 | 14.5 | 19.9 | 24.7 | 13.9 |
+| DS-R1-Distil-Qwen-14B | 10.9 | 15.4 | 13.3 | 14.9 | 11.1 |
+| Llama-4-Scout-17B-16E-Instruct | 6.6 | 10.6 | 6.8 | 7.1 | 7.7 |
+| Llama-3.3-70B-Instruct | 0.0 | 0.9 | 0.0 | 0.3 | 0.4 |
+| Gemini 2.0 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
+| Gemini 2.5 Flash | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
+| Mistral-Large-Instruct-2411 | 0.3 | 0.8 | 0.4 | 1.6 | 0.5 |
+| Claude 3.5 Sonnet | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
+| GPT-4o | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
+Appendix G Hallucinations on Core vs. Peripheral Object Removal
+We qualitatively observe that hallucinations are more frequent when core objects are removed.
+TableĀ 12 presents case-by-case examples where we remove one object at a time and report the corresponding hallucination rates.
+The results show that removing essential objects, such as the pot while cooking, the dishwasher while washing dishes, or the washing machine while doing laundry, leads to high hallucination rates, while removing peripheral objects has little to no effect.
+| Task Name | Corresponding Scene Objects Removed (one at a time) | Model Name | Hallucination Rate |
+| Cook some food | pot, oven, sauce pan | Llama-3-8b-Instruct | pot (54%), others(0%) |
+| Gemma-2-9b-it | pot (52%), others(0%) | ||
+| Wash dishes with dishwasher | dishwasher, dish soap, bowl, plate, drinking glass, | Llama-3-8b-Instruct | dishwasher (100%), plate (38%), others (0%) |
+| fork, cup, spoon, pot, knife | Gemma-2-9b-it | dishwasher (100%), dish soap(18%), plate (39%), others (0%) | |
+| Wash clothes | washing machine, soap, pant, shirt, dress, jacket | Llama-3-8b-Instruct | Washing machine (100%), soap(64%), others (0%) |
+| Gemma-2-9b-it | washing machine (100%), soap(62%), others (0%) |
+Appendix H Hallucination Probing Prompts Design and Distribution
+H.1 Base Prompt Format and Structure
 The base prompts from VirtualHome are more complex, often requiring multi-step reasoning about scene objects and their interactions.
 In contrast, BEHAVIOR prompts closely resemble their corresponding LTL goals, leaving less room for ambiguity and reducing the modelās need to infer or fill in missing information.
-G.2 Prompt Distribution
+H.2 Prompt Distribution
 We report the total number of prompts generated for each hallucination variantāDistractor Injection (DistInj), Task-Relevant Object Removal (TaskObjRem), Synonymous Object Substitution (SynonymSub), and Scene Task Contradiction (SceneTaskCon)āwithin both VirtualHome and BEHAVIOR environments.
@@
 | SceneTaskCon | 100 | |
-Total | 2574 |
-G.3 GPT-4o Prompts
+| Total | 2574 | |
+H.3 GPT-4o Prompts
 FigureĀ 4 and FigureĀ 5 show the prompts used to query GPT-4o to generate our halluciantion probing set.
-Appendix H Model Cards
-TableĀ 12 shows the models cards used in our experiments.
+Appendix I Model Cards
+TableĀ 14 shows the models cards used in our experiments.
 | Model Name | Complete Model ID | Hosting |